OpenStudy (anonymous):

evaluate the integral 3x^2 +1 over x^3 +x

OpenStudy (nikvist):

\[\int\frac{3x^2+1}{x^3+x}dx=\int\frac{d(x^3+x)}{x^3+x}=\ln{|x^3+x|}+C\]